You are here: Foswiki>AdaptiveOptics Web>Logs>Log_20111214 (14 Dec 2011, JuanGuerra)Edit Attach

AO Commissioning on 14 December 2011. LEFT UNIT

AO support: Arcetri team, Juan Carlos Guerra and Michael Wagner (LBTO).

Summary

- Close loop with 400 modes running without disturbance..

- offloading applied to the SX. Working.

- Problems with the CCD39 reading out, noise and it is unknown where is coming from. Checks of the cables is done.

- Problem with the IRTC flux. It seems that we have less flux that we expect.

- We have the network shut down for two days, this stop the work in the AO systems. (This shut down is planned for two days?)

Details.

Positions os the hexapod for W UNIT.

3.5, 0.5,4.37,308,600,0

- The offloading is set in the SX size and it is working. The test has been done at close loop.

- Close loop at bin 1 and the system works fine. Test done with telescope at zenith.

- In close loop and with an artificial star of magnitude 7 we can not see the PSF in the IRTC, the integration time was 4 seconds. The system was rehomed and set up with the sort field with the H filter in and the result is the same (we don't see the psf).

- To check the cables in the W unit, the platform was mounted and we checked the cables and all the connectors. Also a spare AIA board was tested without success. The noise is still there, a more detailed analysis needs to be done. The arcetri team talkj about to remove the box from the unit and to work in the laboratory.

-- JuanGuerra - 14 Dec 2011
Topic revision: r1 - 14 Dec 2011, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback